# RL-AUX: Reinforcement Learning for Auxiliary Task Generation

### v1

@@ Abstract
Auxiliary Learning (AL) is a special case of Multi-task Learning (MTL) in which a network trains on auxiliary tasks to improve performance on its main task. This technique is used to improve generalization and, ultimately, performance on the network’s main task. AL has been demonstrated to improve performance across multiple domains, including navigation, image classification, and natural language processing. One of the main weaknesses of AL is the need for labeled auxiliary tasks, which can require human effort and domain expertise to generate. Meta Learning techniques have been used to solve this issue by learning an additional auxiliary task generation network that can create helpful tasks for the primary network. The most prominent techniques rely on Bi-Level Optimization, which incurs computational cost and increased code complexity. To avoid the need for Bi-Level Optimization, we present an RL-based approach to dynamically create auxiliary tasks. In this framework, an RL agent is tasked with selecting auxiliary labels for every data point in a training set. The agent is rewarded when their selection improves the performance of the main network on the primary task. In this paper, we also experiment with learning optimal strategies for weighing the auxiliary loss per data point. On the 20-Superclass CIFAR100 problem, our RL approach outperforms human-labeled auxiliary tasks and performs as well as one of the most prominent Bi-Level Optimization techniques in the literature. Furthermore, our weight learning approaches significantly outperform all of these benchmarks. For example, our Weight-Aware RL-based approach helps the VGG16 architecture achieve 80.9% test accuracy while the human-labeled auxiliary task setup achieved 75.53%. Ultimately the goal of this work is to (1) prove that RL is a viable approach to dynamically generate auxiliary tasks and (2) demonstrate that per-sample auxiliary task weights can be learned alongside the auxiliary task labels and can achieve strong results.
@@ 1 Introduction
Meta Auxiliary Learning (MAXL) attempts to alleviate the problem of auxiliary task labeling through the procedural generation of an auxiliary task that optimizes the performance on the given primary task (Liu et al., 2019). The MAXL framework works by organizing the inputs of the primary task into hierarchical subclasses for each primary class using an additional label generation network. As such, MAXL is one of several approaches for dynamic label generation that run into the Bi-Level Optimization problem (Navon et al., 2020; Chen et al., 2023). Bi-Level Optimization, which is at the heart of many Meta Learning procedures, arises here as the gradients of the label network are calculated with respect to the performance of the main network on the primary task, resulting in a costly Hessian-inverse vector calculation and increased implementation complexity.
In this work, we will attempt to bypass the need for this complex Bi-Level Optimization altogether by training a Reinforcement Learning (RL) agent to learn the auxiliary task selection. Our approach involves encapsulating the main network’s training loop (for both the primary and auxiliary tasks) within a Reinforcement Learning environment. We then expose the input data points as the environment’s state and the network’s performance on the main task as the reward. The environment’s action space will enable the RL agent to provide auxiliary labels per data point. An agent trained in this environment should therefore learn the auxiliary task labeling that maximizes the performance of the main network on the primary task. We focus our work on commonly used image classification datasets, primarily CIFAR-100. Our experiment results reveal that the RL approach performs as well as the state-of-the-art MAXL approach with less model customization for the Hessian calculation.
@@ 2 Related Work
2.1 Auxiliary and Multi-Task Learning in Vision
Multi-task Learning (MTL) is a well-studied and widely-used Machine Learning method to have a network learn multiple tasks simultaneously (Caruana, 1997; Liebel and Körner, 2018; Zhang and Yang, 2017). Auxiliary Learning (AL) is a special case of MTL, in which there is one primary task of importance and one or more auxiliary tasks that support the performance of the main task (Liu et al., 2019). MTL and AL has been shown to improve target task performance compared to networks that are trained on a single task, particularly in low data contexts (Standley et al., 2019; Zhang and Yang, 2017). There has been extensive work using Multi-task and Auxiliary Learning in the vision domain, such as the auxiliary classifiers in GoogleLeNet (Szegedy et al., 2014), multi-task cascaded convolutional networks (Zhang et al., 2016), state-of-the-art performance on three vision tasks using one convolutional network (Eigen and Fergus, 2014), and many other examples (Strezoski et al., 2023; Kokkinos, 2017; Rasmus et al., 2015). Figure 1(a) represents a common, simple example of an auxiliary task in a classification task setting where the auxiliary task is provided by human labeling.
@@ 2.1 Auxiliary and Multi-Task Learning in Vision
2.2 Task Weighting
The standard loss formulation in Auxiliary Learning is given by:
@@ 2.2 Task Weighting
2.3 Task Generation
MAXL started the label network paradigm for dynamic task generation in the Auxiliary Learning space using Bi-Level Optimization (Liu et al., 2019). The MAXL framework trains a label network to create an auxiliary task that optimizes the main network’s performance on the target primary task. Figure 1(b) shows how the setup of a label network is used to generate the auxiliary task. Bi-Level Optimization approaches attempt to find the optimal primary network weights $\theta^{*}$ and auxiliary labeling network weights $\phi^{*}$ that satisfy:
@@ 2.3 Task Generation
2.4 Combining Auxiliary Learning and Reinforcement Learning
The majority of work combining Reinforcement Learning with Auxiliary Learning involves constructing auxiliary tasks for RL agents to perform better on their assigned task (Jaderberg et al., 2016; Riedmiller et al., 2018; Narvekar et al., 2017).
@@ 3 Problem Formulation
3.1 Reinforcement Learning Approach
The policy $\pi_{\phi}$ will be iteratively updated through an RL approach. To do this, we encapsulate the learning of the main network within an environment. Figure 2 provides an overview of how the training is set up.
@@ 3.1 Reinforcement Learning Approach
3.2 Label Hierarchy Optimization
MAXL employs hierarchy-constrained label generation through the use of a Masked Soft-max (Liu et al., 2019). In other words, each class in the primary task will have some fixed number of subclasses in the auxiliary task.
@@ 3.2 Label Hierarchy Optimization
3.3 Weight-Adjusted Approaches
After initial successes with the RL task generation approach mentioned, we turned our attention to learning sample-level weight adjustment in both the Bi-Level Optimization and RL approaches. Our per-sample loss from Equation 2 is adjusted slightly, so that the weighting is sample-specific:
@@ 3.3 Weight-Adjusted Approaches
For the Reinforcement Learning approach, we added another component to the action space, to enable the agent to select a weight for the given data point. The Weight-Adjusted Reinforcement Learning architecture is presented in Figure 9(b).
For the Bi-Level Optimization approach, we took MAXL’s open-source implementation and added a weight-output head to the labeling network such that it also outputs a weight parameter for a given sample. This new Weight-Adjusted architecture, WA-MAXL, is presented in Figure 9(a).
(a) Weight-Aware Label Network Auxiliary Task Setup
(b) Weight-Aware Reinforcement Learning Auxiliary Task Setup
Figure 3: Weight-Aware Auxiliary Task Approaches
@@ 4 Methodology
4.1 RL Environment Implementation
We implemented the main network environment (see Figure 2) using OpenAI’s Gymnasium framework (Brockman et al., 2016). Our custom environment is configured to be agnostic of the primary task, type of main network, type of RL agent, system hardware, optimizer/scheduler, etc. These parameters can be set during instantiation, making our environment a robust platform to prototype Auxiliary Task training on different architectures and datasets rapidly.
4.1.1 Agent Interaction
Our environment is written to operate with an observation space that provides a single data point (image). The action it expects back is the RL agent’s labeling of the last provided image. Therefore, on start-up, the environment samples a single batch from the training dataloader and provides a single data point (image) from the batch to the RL agent as the observation. The other data points in the batch are held for use as future observations. The RL agent then selects the action (label) for the observation and provides it to the environment.
@@ 4.1.1 Agent Interaction
4.1.2 Training Modes
In addition to the standard OpenAI Gymnasium interface, our custom environment has additional features to support our unique training. This includes training evaluation functions, main network persisting, metric capturing, and so on.
@@ 4.1.2 Training Modes
4.2 Weight Selection
The current version of Stable-Baselines3’s $MultiInputPolicy$ does not support combining discrete label actions ($MultiDiscrete$) and continuous scalar actions ($Box$) simultaneously (Raffin et al., 2021). As such, when implementing the weight-aware version of the RL approach, we treat weight selection in the action space as a 21-class classification problem. Each class represented an evenly spaced value between 0 and 1, as in $\{0,0.05,0.1,0.15,...,0.95,1\}$. The selected unscaled weight $w_{u}$, is then used to get the true selected weight as:
@@ 4.2 Weight Selection
4.3 Networks
4.4 Main Network
4.5 Agent Network
where
and $A_{t}$ denotes the advantage.
Concretely, our implementation is composed of 3 custom components:
4.6 Datasets
Cifar-10 and Cifar-100 have become ubiquitous benchmark datasets in the Auxiliary Learning space, both of which contain 60,000 images spread across 10 and 100 classes, respectively (Grégoire et al., 2023; Chen et al., 2023; Fan et al., 2018; Cubuk et al., 2018; Liu et al., 2019; Navon et al., 2020; Krizhevsky, 2009).
Cifar-100 provides a 20 superclass hierarchy. Each of the 100 classes in the dataset is mapped into one of the 20 Superclasses such that each superclass is composed of 5 of the 100 classes (Krizhevsky, 2009). Table 1 in the Appendix shows how the Superclasses are constructed. This is a very useful human-labeled auxiliary task benchmark, against which we can compare our proposed methodology.
SVHN (Street View House Numbers) is also a prominent test dataset in the literature (Netzer et al., 2011). It is composed of 600,000 real-world images of address digits with 10 classes (0-9).
Our experiments will explore how the main network performance changes on these datasets as we introduce RL-created auxiliary tasks.
4.7 Training
The training procedure was standardized in all experiments, unless otherwise noted. The training strategy involved alternating the training between one epoch of RL Agent training followed by one epoch of main network training. The two supported training modes mentioned in Section 4.1.2 facilitated this alternating training strategy. The environment was reset (dataloader was reshuffled, canonical main network was loaded, global state cleared, etc.) between batches. Figure 4 visualizes this workflow.
Figure 4: Alternating Training Approach
[…]
@@ 5 Results and Analysis
5.1 Reinforcement Learning vs MAXL
To judge the performance benefits of our RL-based Auxiliary Task generation approach, we compare it to: (1) a network trained with a MAXL-generated Auxiliary Task, (2) a Human-Labeled Auxiliary Task (see Section 4.6), and (3) a network trained without an auxiliary task. We will test these approaches on the CIFAR-100 20-Superclass dataset. In this experiment, the auxiliary tasks being tested are equally weighted to the primary task.
@@ 5.1 Reinforcement Learning vs MAXL
In Section 5.3, we will report results of the RL-based approach on additional datasets.
@@
5.2 Weight Adjustment Ablation
Next we wanted to observe how tuning the task-level weight hyperparameter $\lambda$ affects the performance of the main network on the primary task. To do this, we retrained the RL Agent and the main network on the Cifar100 20-Superclass task with $\lambda=0.25,0.5,1,2,4$ and observed the results.
@@ 5.2 Weight Adjustment Ablation
5.3 Weight-Aware Approach Experiment
Now we wish to test the performance of our weight-aware approaches on the 20-Superclass CIFAR100 task. As is clear in Figure 7, our Weight-Aware MAXL (WA-MAXL) and Weight-Aware RL approach (WA-RL) provide significant performance improvement over the other methods. On the 20-Superclass Cifar100 task, WA-MAXL is the highest performer with 81.3% accuracy, followed closely by WA-RL with 80.9% accuracy. The non-weight-aware approaches hover in the 76% range, while the single task network is just below 72%

### v4

# Generating Auxiliary Tasks with Reinforcement Learning

@@ Abstract
Auxiliary Learning (AL) is a form of multi-task learning in which a model trains on auxiliary tasks to boost performance on a primary objective. While AL has improved generalization across domains such as navigation, image classification, and NLP, it often depends on human-labeled auxiliary tasks that are costly to design and require domain expertise. Meta-learning approaches mitigate this by learning to generate auxiliary tasks, but typically rely on gradient based bi-level optimization, adding substantial computational and implementation overhead. We propose RL-AUX, a reinforcement-learning (RL) framework that dynamically creates auxiliary tasks by assigning auxiliary labels to each training example, rewarding the agent whenever its selections improve the performance on the primary task. We also explore learning per-example weights for the auxiliary loss. On CIFAR-100 grouped into 20 superclasses, our RL method outperforms human-labeled auxiliary tasks and matches the performance of a prominent bi-level optimization baseline. We present similarly strong results on other classification datasets. These results suggest RL is a viable path to generating effective auxiliary tasks.
@@ 1 Introduction
Meta Auxiliary Learning (MAXL) attempts to alleviate the problem of auxiliary task labeling through the procedural generation of an auxiliary task that optimizes the performance on the given primary task (Liu et al., 2019). The MAXL framework works by organizing the inputs of the primary task into hierarchical subclasses for each primary class using an additional label generation network. As such, MAXL is one of several approaches for dynamic label generation that run into the Bi-Level Optimization problem (Navon et al., 2020; Chen et al., 2023). Bi-Level Optimization, which is at the heart of many Meta Learning procedures, arises here as the gradients of the label network are calculated with respect to the performance of the main network on the primary task, resulting in a Hessian-inverse vector calculation and increased implementation complexity.
In this work, we will attempt to bypass the need for this complex Bi-Level Optimization altogether by training a Reinforcement Learning (RL) agent to learn the auxiliary task selection (RL-AUX). Our approach involves encapsulating the main network’s training loop (for both the primary and auxiliary tasks) within a Reinforcement Learning environment. We then expose the input data points as the environment’s state and the network’s performance on the main task as the reward. The environment’s action space will enable the RL agent to provide auxiliary labels per data point. An agent trained in this environment should therefore learn the auxiliary task labeling that maximizes the performance of the main network on the primary task. We focus our work on commonly used image classification datasets, primarily CIFAR-100. Our experiment results reveal that the RL approach performs as well as the state-of-the-art MAXL approach with less model customization for the Hessian calculation.
@@ 2 Related Work
Auxiliary and Multi-Task Learning in Vision Multi-task Learning (MTL) is a well-studied and widely-used Machine Learning method to have a network learn multiple tasks simultaneously (Caruana, 1997; Liebel and Körner, 2018; Zhang and Yang, 2017). Auxiliary Learning (AL) is a special case of MTL, in which there is one primary task of importance and one or more auxiliary tasks that support the performance of the main task (Liu et al., 2019). MTL and AL has been shown to improve target task performance compared to networks that are trained on a single task, particularly in low data contexts (Standley et al., 2019; Zhang and Yang, 2017). There has been extensive work using Multi-task and Auxiliary Learning in the vision domain, such as the auxiliary classifiers in GoogleLeNet (Szegedy et al., 2014), multi-task cascaded convolutional networks (Zhang et al., 2016), state-of-the-art performance on three vision tasks using one convolutional network (Eigen and Fergus, 2014), and many other examples (Strezoski et al., 2023; Kokkinos, 2017; Rasmus et al., 2015). Figure 1(a) represents a common, simple example of an auxiliary task in a classification task setting where the auxiliary task is provided by human labeling.
@@
Task Weighting The standard loss formulation in Auxiliary Learning is given by:
@@
Task Generation MAXL started the label network paradigm for dynamic task generation in the Auxiliary Learning space using Bi-Level Optimization (Liu et al., 2019). The MAXL framework trains a label network to create an auxiliary task that optimizes the main network’s performance on the target primary task. Figure 1(b) shows how the setup of a label network is used to generate the auxiliary task. Bi-Level Optimization approaches attempt to find the optimal primary network weights $\theta^{*}$ and auxiliary labeling network weights $\phi^{*}$ that satisfy:
@@
Combining Auxiliary Learning and Reinforcement Learning The majority of work combining Reinforcement Learning with Auxiliary Learning involves constructing auxiliary tasks for RL agents to perform better on their assigned task (Jaderberg et al., 2016; Riedmiller et al., 2018; Narvekar et al., 2017).
@@ 3 Problem Formulation
Reinforcement Learning Approach The policy $\pi_{\phi}$ will be iteratively updated through an RL approach. To do this, we encapsulate the learning of the main network within an environment. Figure 2 provides an overview of how the training is set up.
@@
Label Hierarchy Optimization MAXL employs hierarchy-constrained label generation through the use of a Masked Soft-max (Liu et al., 2019). In other words, each class in the primary task will have some fixed number of subclasses in the auxiliary task.
@@
Weight-Adjusted Approaches After initial successes with the RL task generation approach mentioned, we turned our attention to learning sample-level weight adjustment in both the Bi-Level Optimization and RL approaches. Our per-sample loss from Equation 2 is adjusted slightly, so that the weighting is sample-specific:
@@
For the Reinforcement Learning approach, we added another component to the action space, to enable the agent to select a weight for the given data point. The Weight-Adjusted Reinforcement Learning architecture is presented in Figure 3.
For the Bi-Level Optimization approach, we took MAXL’s open-source implementation and added a weight-output head to the labeling network such that it also outputs a weight parameter for a given sample.
Figure 3: Weight-Aware Reinforcement Learning Auxiliary Task Setup
@@ 4 Methodology
RL Environment Implementation We implemented the main network environment (see Figure 2) using OpenAI’s Gymnasium framework (Brockman et al., 2016). Our custom environment is configured to be agnostic of the primary task, type of main network, type of RL agent, system hardware, optimizer/scheduler, etc. These parameters can be set during instantiation, making our environment a robust platform to prototype Auxiliary Task training on different architectures and datasets rapidly.
Agent Interaction Our environment is written to operate with an observation space that provides a single data point (image). The action it expects back is the RL agent’s labeling of the last provided image. Therefore, on start-up, the environment samples a single batch from the training dataloader and provides a single data point (image) from the batch to the RL agent as the observation. The other data points in the batch are held for use as future observations. The RL agent then selects the action (label) for the observation and provides it to the environment.
@@
Training Modes In addition to the standard OpenAI Gymnasium interface, our custom environment has additional features to support our unique training. This includes training evaluation functions, main network persisting, metric capturing, and so on.
@@
Weight Selection The current version of Stable-Baselines3’s $MultiInputPolicy$ does not support combining discrete label actions ($MultiDiscrete$) and continuous scalar actions ($Box$) simultaneously (Raffin et al., 2021). As such, when implementing the weight-aware version of the RL approach, we treat weight selection in the action space as a 21-class classification problem. Each class represented an evenly spaced value between 0 and 1, as in $\{0,0.05,0.1,0.15,...,0.95,1\}$. The selected unscaled weight $w_{u}$, is then used to get the true selected weight as:
@@
Datasets Cifar-10 and Cifar-100 have become ubiquitous benchmark datasets in the Auxiliary Learning space, both of which contain 60,000 images spread across 10 and 100 classes, respectively (Grégoire et al., 2023; Chen et al., 2023; Fan et al., 2018; Cubuk et al., 2018; Liu et al., 2019; Navon et al., 2020; Krizhevsky, 2009). Cifar-100 provides a 20 superclass hierarchy. Each of the 100 classes in the dataset is mapped into one of the 20 Superclasses such that each superclass is composed of 5 of the 100 classes (Krizhevsky, 2009). Table 1 in the Appendix shows how the Superclasses are constructed. This is a very useful human-labeled auxiliary task benchmark, against which we can compare our proposed methodology. SVHN (Street View House Numbers) is also a prominent test dataset in the literature (Netzer et al., 2011). It is composed of 600,000 real-world images of address digits with 10 classes (0-9). Our experiments will explore how the main network performance changes on these datasets as we introduce RL-created auxiliary tasks. For further training details, see Appendix B. For network details, see Appendix C.
@@ 5 Results and Analysis
Reinforcement Learning vs MAXL To judge the performance benefits of our RL-based Auxiliary Task generation approach, we compare it to: (1) a network trained with a MAXL-generated Auxiliary Task, (2) a Human-Labeled Auxiliary Task (see Section 4), and (3) a network trained without an auxiliary task. We will test these approaches on the CIFAR-100 20-Superclass dataset. In this experiment, the auxiliary tasks being tested are equally weighted to the primary task.
@@
In Section 5, we will report results of the RL-based approach on additional datasets.
@@
Weight Adjustment Ablation Next we wanted to observe how tuning the task-level weight hyperparameter $\lambda$ affects the performance of the main network on the primary task. To do this, we retrained the RL Agent and the main network on the Cifar100 20-Superclass task with $\lambda=0.25,0.5,1,2,4$ and observed the results.
@@
Weight-Aware Approach Experiment Now we wish to test the performance of our weight-aware approaches on the 20-Superclass CIFAR100 task. As is clear in Figure 6, our Weight-Aware MAXL (WA-MAXL) and Weight-Aware RL approach (WA-RL) provide significant performance improvement over the other methods. On the 20-Superclass Cifar100 task, WA-MAXL is the highest performer with 81.3% accuracy, followed closely by WA-RL with 80.9% accuracy. The non-weight-aware approaches hover in the 76% range, while the single task network is just below 72%
@@ A.3 Precision, Recall and F1 Score
Appendix B Training
The training procedure was standardized in all experiments, unless otherwise noted. The training strategy involved alternating the training between one epoch of RL Agent training followed by one epoch of main network training. The two supported training modes mentioned in Section 4 facilitated this alternating training strategy. The environment was reset (dataloader was reshuffled, canonical main network was loaded, global state cleared, etc.) between batches. Figure 13 visualizes this workflow.
Figure 13: Alternating Training Approach
[…]
Appendix C Networks
C.1 Main Network
C.2 Agent Network
where
and $A_{t}$ denotes the advantage.
Concretely, our implementation is composed of 3 custom components: